# Sample from What You See: Visuomotor Policy Learning via Diffusion Bridge with Observation-Embedded Stochastic Differential Equation

### v1

@@ Abstract
Imitation learning with diffusion models has advanced robotic control by capturing multi-modal action distributions. However, existing approaches typically treat observations as high-level conditioning inputs to the denoising network, rather than integrating them into the stochastic dynamics of the diffusion process itself. As a result, sampling must begin from random Gaussian noise, weakening the coupling between perception and control and often yielding suboptimal performance. We introduce BridgePolicy, a generative visuomotor policy that explicitly embeds observations within the stochastic differential equation via a diffusion-bridge formulation. By constructing an observation-informed trajectory, BridgePolicy enables sampling to start from a rich, informative prior rather than random noise, substantially improving precision and reliability in control. A key challenge is that classical diffusion bridges connect distributions with matched dimensionality, whereas robotic observations are heterogeneous and multi-modal and do not naturally align with the action space. To address this, we design a multi-modal fusion module and a semantic aligner that unify visual and state inputs and align observation and action representations, making the bridge applicable to heterogeneous robot data. Extensive experiments across 52 simulation tasks on three benchmarks and five real-world tasks demonstrate that BridgePolicy consistently outperforms state-of-the-art generative policies.
@@ 1 Introduction
Figure 1: Comparison of Diffusion Policy and our BridgePolicy. The observation modeling way of BridgePolicy allows its sampling of BridgePolicy can start from an rich and meaningful prior instead of the random noise in standard Diffusion Policy.
@@
**footnotetext: Corresponding author.
Imitation learning (IL)  [osa2018algorithmic] is a widely adopted learning paradigm in robotic learning [li2024robust, shafiullah2023bringing, bin2020learning, ze2023gnfactor, seo2023deep, fu2024mobile], where a robot is provided with a set of expert demonstrations and learns to mimic the provided demonstrations to perform the tasks effectively. Recently, generative models such as diffusion model [ho2020denoising, score-based, classifier, classifier-free] and flow matching [lipman2024flow, flow_for_generative, flow_fast, SI] gains its prominence in IL owing to their capacity to fit multi-modal distributions and learn temporal dependency [chi2023diffusion, ze2023visual, zhang2025flowpolicy]. These methods share a similar principle: they perturb action trajectories into random noise via a forward process defined by a Stochastic or Ordinary Differential Equation (SDE/ODE) and then train a neural network conditioned on observations to reverse this process, iteratively transforming noise into executable actions.
@@
With this paradigm, Diffusion Policy [chi2023diffusion] and 3D Diffusion Policy [ze2023visual], known as DP and DP3, respectively, employ an SDE-defined forward process and train the neural network conditioned on visual inputs and robot states to control the denoising process to sample actions from the random noise. FlowPolicy [zhang2025flowpolicy], in contrast, employs an ODE-defined forward and reverse process, which reduces the stochasticity during training and inference. Despite their success, current generative policies largely treat observations as high-level conditioning signals to the denoising network [chi2023diffusion, ze2023visual, zhang2025flowpolicy], rather than integrating them into the dynamics of the forward process. This underutilization forces sampling to begin from uninformative Gaussian noise, weakening the coupling between perception and control and often yielding suboptimal performance.
@@
Diffusion bridge has demonstrated considerable success in the image generation and translation domain [yue2023image, li2023bbdm, luo2023image, de2021diffusion, zhou2023denoising], where it directly modifies the forward process such that the endpoint distribution naturally aligns with the desired conditioned distribution. As shown in Figure 1, with this mathematically exact formulation of modeling informative observations in the forward process rather than solely treating the observations as the external condition, the reverse process can naturally start from the observations, a more informative prior than the uninformative random noise, thereby improving the precision of the generated actions. Building upon these insights, we propose that diffusion bridge can also serve as a more effective framework for visuomotor policy learning. Specifically, rather than adopting the conventional diffusion model, we formulate policy learning as the problem of learning a diffusion bridge, where observations are explicitly modeled in to the diffusion SDE trajectory itself through the framework of diffusion bridge [zhu2025unidb, pan2025unidb] instead of merely treating them as conditions, and the reverse process would be able to sample actions starting from a more informative prior of the observations instead of the uninformative random Gaussian noise. This deeper integration allows the policy to more effectively exploit observations, leading to more precise and reliable control.
However, formulating policy learning as the diffusion bridge raises two key challenges. First, robotic observations and actions are inherently multi-modal and heterogeneous, violating the standard diffusion-bridge assumption that the connected endpoint distributions share the same dimensionality. Second, observations often include proprioceptive states, RGB-D vision, and language instructions, which do not admit a simple one-to-one mapping to the action space required by classical bridge formulations.
@@
We introduce BridgePolicy, a generative visuomotor policy that directly samples actions from observation-informed priors rather than random noise. We construct a diffusion bridge that embeds observations within the diffusion SDE trajectory, enabling the model to fully exploit sensory inputs instead of using them solely as external conditions. To resolve the modality and shape mismatches, we design two components: (i) a multi-modal fusion module that consolidates heterogeneous inputs (e.g., vision, proprioception, language) into a unified observation representation, and (ii) a semantic aligner that maps this fused representation into an action-aligned latent space, ensuring compatible endpoints for the bridge. Together, these modules allow the diffusion bridge to leverage heterogeneous data sources for policy learning. Our contributions are summarized as follows:
@@
  • To make diffusion bridges applicable to heterogeneous robotic data, we introduce a multi-modal fusion module and a semantic aligner that unify visual and state observations and align their representations with the action space. This resolves modality and shape mismatches, enabling effective use of multi-modal observations for policy learning.
@@ 2.1 Generative Models in Robotic Learning
Recent years have witnessed the rise of diffusion models and flow matching as the state-of-the-art generative paradigms, largely driven by their remarkable success in the image generation and reinforcement learning domain [ho2020denoising, song2020denoising, rombach2022high, lipman2022flow, afforddp, QVPO, ding2025exploring, ding2025genpo], etc. Due to their capacity of multi-modal expressiveness, they are widely used in areas such as reinforcement learning, imitation learning, and motion planning [janner2022planning, reuss2023goal, chi2023diffusion, pearce2023imitating, sridhar2023memory, xian2023chaineddiffuser, prasad2024consistency, saha2024edmp]. In robotics, both diffusion models and flow matching have been utilized as policy learning frameworks, where action chunks are generated from a random noise under the condition of 2D or 3D observations through the SDE or ODE-defined generative processes, respectively [chi2023diffusion, ze20243d, zhang2025flowpolicy, hu2024adaflow]. In contrast, our method is pioneering in modeling the observations in the diffusion SDE trajectory via SOC. The sampling therefore starts from a more informative prior, allowing the policy to more effectively exploit observations and generate more precise and reliable behaviors.
@@ 2.2 Diffusion Bridges for Generative Modeling
Diffusion bridges enable the transition between two arbitrary distributions without the need to initiate the diffusion process from the random noise, which provides a more principled and mathematically exact generation paradigm for one-to-one distribution transformation. On the one hand, Diffusion Schrödinger Bridges [liu20232, shi2023diffusionschrodingerbridgematching, debortoli2023diffusionschrodingerbridgeapplications, somnath2024aligneddiffusionschrodingerbridges, chen2021likelihood, deng2024variational, wang2025implicit] aim to determine a stochastic process by solving an entropy-regularized optimal transport problem between two distributions. However, its high computational complexity, particularly pronounced in high-dimensional settings or constraints, poses significant challenges for direct optimization. Other works [zhou2023denoising, yue2023image] incorporate Doob’s $h$-transform into the forward SDE process, delivering remarkable performance on image restoration benchmarks. UniDB [zhu2025unidb] reformulates diffusion bridges through an SOC-based optimization problem, proving that Doob’s $h$-transform is a special case of SOC theory, thereby unifying and generalizing existing $h$-transform-based diffusion bridge methods. In this work, we leverage the ability of the diffusion bridge to connect heterogeneous distributions, specifically the multi-modal observation distribution and the action distribution of an expert policy, and harness this mechanism for policy learning.
@@ 3.1 BridgePolicy
Existing policy learning frameworks [chi2023diffusion, ze20243d, zhang2025flowpolicy, hu2024adaflow] aim to learn a policy $\pi:\boldsymbol{O}\rightarrow\boldsymbol{A}$, which predict the actions $\boldsymbol{a}\in\boldsymbol{A}$ given the observations $\boldsymbol{o}\in\boldsymbol{O}$ which include 3D point cloud as the visual input and robot state and the actions are chunked into a short sequence of a trajectory finishing the task. Previous works such as DP3 [ze20243d] and FlowPolicy [zhang2025flowpolicy] formulate the policy as a conditional generative model, where actions are generated from random noises under the condition of observations $\boldsymbol{o}$ [chi2023diffusion, ze20243d, zhang2025flowpolicy]. They merely treated the informative observations as conditional signals to guide the neural network during denoising.
Forward Process of BridgePolicy. Motivated by the recent advances of Diffusion Bridge [zhou2023denoising, yue2023image, zhu2025unidb], where conditions can be explicitly modeled in the forward diffusion SDE trajectory, we directly integrate the observations into the diffusion process to generate a more controllable and precise robot action. Here, our BridgePolicy adopt a unified diffusion bridge framework [zhu2025unidb], formulated under the Stochastic Optimal Control (SOC) theory, as the core learning paradigm. To be consistent with the notations of previous diffusion for visuomotor policy learning [chi2023diffusion], we denote $\boldsymbol{a}_{t}$ as the interpolated robot action state at time $t$. Assuming that the action $\boldsymbol{a}$ and the observation $\boldsymbol{o}$ share the same dimension, we define the endpoints of our forward diffusion bridge process (2) as follows:
@@
  • The terminal state $\boldsymbol{a}_{T}$: we set the terminal state to be the observation, i.e., $\boldsymbol{a}_{T}=\boldsymbol{o}$. It is a pivotal departure from standard diffusion-based or flow-based policy where $\boldsymbol{a}_{T}\sim\mathcal{N}(0,I)$ is typically random Gaussian noise.
According to UniDB [zhu2025unidb], we can construct the forward process of diffusion bridge as the following SOC optimization problem:
@@
  $\boldsymbol{z}_{obs}=\mathbf{softmax}({\boldsymbol{z}_{s}\cdot\boldsymbol{z}_{s}^{T}}/{\sqrt{d_{s}}})\boldsymbol{z}_{pc}$
@@
$\begin{gathered}\min_{\mathbf{u}_{t,\gamma}}\ \mathbb{E}\left[\int_{0}^{T}\frac{1}{2}\left\|\mathbf{u}_{t,\gamma}\right\|_{2}^{2}dt+\frac{\gamma}{2}\left\|\boldsymbol{a}_{T}^{u}-\boldsymbol{a}_{T}\right\|_{2}^{2}\right]\\ \text{s.t.}\ \mathrm{d}\boldsymbol{a}_{t}=\Big[\theta_{t}(\boldsymbol{a}_{T}-\boldsymbol{a}_{t})+g_{t}\mathbf{u}_{t,\gamma}\Big]\mathrm{d}t+g_{t}\mathrm{d}\boldsymbol{w}_{t},\ \boldsymbol{a}_{0}^{u}=\boldsymbol{a}_{0},\end{gathered}$ (1)
@@
where we follow the $l_{1}$ norm in UniDB [zhu2025unidb] regarding the order of the norm, and provide the ablation study in the Supplementary Materials. Consequently, BridgePolicy naturally integrates the informative observations $\boldsymbol{a}_{T}=\boldsymbol{o}$ both in the SDE trajectories and the neural networks and aims to learn a stochastic trajectory that transforms the observation $\boldsymbol{o}$ into a plausible action $\boldsymbol{a}$.
@@
 Initialize $\boldsymbol{a}_{T}=\boldsymbol{z}_{obs}=\mathbf{softmax}({\boldsymbol{z}_{s}\cdot\boldsymbol{z}_{s}^{T}}/{\sqrt{d_{s}}})\boldsymbol{z}_{pc}$
@@
Return Actions $\tilde{\boldsymbol{a}}$
Fast Sampling Algorithm. After we obtain the optimal model $\boldsymbol{a}^{*}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$, the decision making module of our BridgePolicy can be implemented as an iterative generation procedure starting from the given observation $\boldsymbol{a}_{T}=\boldsymbol{o}$. As for specially designed solvers, UniDB++ [pan2025unidb] provides a training-free acceleration algorithm with the data prediction $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$ which directly estimates the fixed and smooth target $\boldsymbol{a}_{0}$. With some notations $\rho_{t}=e^{\bar{\theta}_{t}}(1-e^{-2\bar{\theta}_{t}})$, $\kappa_{t,\gamma}=e^{\bar{\theta}_{t:T}}((\gamma\lambda^{2})^{-1}+1-e^{-2\bar{\theta}_{t:T}})$, $c_{1}=(\gamma\lambda^{2})^{-1}e^{2\bar{\theta}_{T}}$, $c_{2}=e^{2\bar{\theta}_{T}}-1$, $D=2c_{1}c_{2}/(c_{1}+c_{2})^{3}$, $E=c_{2}^{2}/(c_{1}+c_{2})^{2}$, and $F=c_{1}^{2}/(c_{1}+c_{2})^{2}$ in UniDB++, the closed-form updating rule from time step $t$ to $s$ is formed as
$\begin{gathered}\begin{aligned} \boldsymbol{a}_{t}=&\frac{\kappa_{t,\gamma}\kappa_{s}\rho_{t}}{\kappa_{s,\gamma}\kappa_{t}\rho_{s}}\boldsymbol{a}_{s}+(1-\frac{\kappa_{t,\gamma}\kappa_{s}\rho_{t}}{\kappa_{s,\gamma}\kappa_{t}\rho_{s}}+\frac{\kappa_{t,\gamma}\rho_{t}\kappa_{s}}{\kappa_{0,\gamma}\kappa_{t}\rho_{s}}-\frac{\kappa_{t,\gamma}}{\kappa_{0,\gamma}})\boldsymbol{a}_{T}\\ &+\left(\frac{\kappa_{t,\gamma}}{\kappa_{0,\gamma}}-\frac{\kappa_{t,\gamma}\rho_{t}\kappa_{s}}{\kappa_{0,\gamma}\kappa_{t}\rho_{s}}\right)\boldsymbol{a}_{\theta}(\boldsymbol{a}_{s},\boldsymbol{a}_{T},s)+\delta^{d}_{s:t}\boldsymbol{\epsilon},\end{aligned}\\ \begin{aligned} (\delta^{d}_{s:t})^{2}&=\frac{\lambda^{2}\kappa^{2}_{t,\gamma}\rho^{2}_{t}}{\kappa^{2}_{t}}\Bigg[E\frac{e^{2\bar{\theta}_{s}}-e^{2\bar{\theta}_{t}}}{(e^{2\bar{\theta}_{t}}-1)(e^{2\bar{\theta}_{s}}-1)}\\ &+D\log\frac{\kappa_{s,\gamma}\rho_{t}}{\kappa_{t,\gamma}\rho_{s}}-F(\frac{e^{-\bar{\theta}_{T}-\bar{\theta}_{t}}}{\kappa_{t,\gamma}}-\frac{e^{-\bar{\theta}_{T}-\bar{\theta}_{s}}}{\kappa_{s,\gamma}})\Bigg],\end{aligned}\end{gathered}$ (4)

### v2

@@ Abstract
Imitation learning with diffusion models has advanced robotic control by capturing the multi-modal action distributions. However, existing methods typically treat observations only as high-level conditions to the denoising network, rather than integrating them into the stochastic dynamics of the diffusion process itself. As a result, the sampling is forced to begin from random noise, weakening the coupling between perception and control and often yielding suboptimal performance. We propose BridgePolicy, a generative visuomotor policy that directly integrates observations into the stochastic dynamics via a diffusion-bridge formulation. By constructing an observation-informed trajectory, BridgePolicy enables sampling to start from a rich and informative prior rather than random noise, substantially improving precision and reliability in control. A key difficulty is that diffusion bridge normally connects distributions of matched dimensionality, while robotic observations are heterogeneous and not naturally aligned with actions. To overcome this, we introduce a multi-modal fusion module and a semantic aligner to unify the visual and state inputs and align the observations with action representations, making diffusion bridge applicable to heterogeneous robot data. Extensive experiments across 52 simulation tasks on three benchmarks and 5 real-world tasks demonstrate that BridgePolicy consistently outperforms state-of-the-art generative policies.
Machine Learning, ICML
@@ 1 Introduction
Imitation learning (Osa et al., 2018) is a widely adopted learning paradigm in robotic learning (Li et al., 2024; Shafiullah et al., 2023; Ze et al., 2023b; Seo et al., 2023; Fu et al., 2024), where a robot is provided with a set of expert demonstrations and learns to mimic the provided demonstrations to perform the tasks effectively. Recently, generative models such as diffusion model (Ho et al., 2020; Song et al., 2020b; Dhariwal and Nichol, 2021; Ho and Salimans, 2022) and flow matching (Lipman et al., 2024; Liu et al., 2022; Albergo et al., 2023) gains its prominence owing to their capacity to capture multi-modal distributions and learn temporal dependency (Chi et al., 2023; Ze et al., 2023a; Zhang et al., 2025). These methods share a similar principle that perturb action chunks into random noise via a forward process defined by a Stochastic or Ordinary Differential Equation (SDE/ODE) and then train a neural network conditioned on observations to reverse this process, iteratively transforming noise into executable actions.
Figure 1: Comparison of Diffusion Policy and our BridgePolicy. The observation modeling way of BridgePolicy allows its sampling of BridgePolicy can start from an informative prior instead of the random noise in Diffusion Policy.
@@
With this paradigm, Diffusion Policy (Chi et al., 2023) and 3D Diffusion Policy (Ze et al., 2023a), known as DP and DP3, employ an SDE-defined forward process and train the neural network conditioned on visual inputs and robot states to control the denoising process to sample actions from the random noise. Similarly, FlowPolicy (Zhang et al., 2025) employs an ODE-defined forward and reverse process, which reduces the stochasticity during training and inference. Despite their success, current generative policies largely treat observations as high-level conditioning signals to the denoising network (Chi et al., 2023; Ze et al., 2023a; Zhang et al., 2025), rather than integrating them into the dynamics of the forward process. This underutilization forces sampling to begin from an uninformative random noise, weakening the coupling between perception and control and often yielding suboptimal performance.
@@
Diffusion bridge has demonstrated considerable success in image restoration and translation (De Bortoli et al., 2021a; Yue et al., 2023; Li et al., 2023; Luo et al., 2023; Zhou et al., 2023), where it modifies the forward process such that the endpoint distribution naturally aligns with the desired conditioned distribution in standard diffusion. As shown in Figure 1, with this mathematically exact formulation of modeling informative observations in the forward process rather than solely treating the observations as the external condition, the reverse process can naturally start from the observations, a more informative prior instead of the uninformative random noise, thereby improving the precision of the generated actions. Building upon these insights, we propose that diffusion bridge can also serve as a more effective framework for visuomotor policy learning. Specifically, rather than adopting the conventional diffusion model, we formulate policy learning as the problem of learning a diffusion bridge, where observations are explicitly modeled in to the diffusion SDE trajectory itself through the framework of diffusion bridge (Zhu et al., 2025; Pan et al., 2025) instead of merely treating them as conditions, and the reverse process would be able to sample actions starting from a more informative prior of the observations instead of the uninformative random Gaussian noise. This deeper integration allows the policy to more effectively exploit observations, leading to more precise and reliable control.
@@
However, formulating policy learning as the diffusion bridge brings two difficulties. First, robotic observations and actions are inherently multi-modal and heterogeneous, violating the standard diffusion-bridge assumption that the connected endpoint distributions share the same dimensionality. Second, observations often include proprioceptive states, RGB-D vision, and language instructions, which do not admit a simple one-to-one mapping to the action space required by classical bridge formulations. Therefore, unifying the representation of multi-modal observations and aligning the observation and action shapes constitute the critical challenges in making diffusion bridges applicable to policy learning. We introduce BridgePolicy, a generative visuomotor policy that directly samples actions from observation-informed priors rather than random noise. We construct a diffusion bridge that embeds observations within the diffusion SDE trajectory, enabling the model to fully exploit sensory inputs instead of using them solely as external conditions. To resolve the modality and shape mismatches, we design two components: (i) a multi-modal fusion module that consolidates heterogeneous inputs (e.g., vision, proprioception) into a unified observation representation, and (ii) a semantic aligner that maps this fused representation into an action-aligned latent space, ensuring compatible endpoints for the bridge. Together, these modules allow the diffusion bridge to leverage heterogeneous data sources for policy learning. Our contributions are summarized as follows:
@@
  • To make diffusion bridge applicable to heterogeneous robotic data, we introduce a multi-modal fusion module and a semantic aligner that unify visual and state observations and align their representations with the action space. This resolves modality and shape mismatches, enabling effective use of multi-modal observations.
@@ 2.1 Generative Models in Robotic Learning
Diffusion models become state-of-the-art generative models, driven by their success in image generation (Song et al., 2020a; Rombach et al., 2022; Lipman et al., 2022). Due to the capacity of multi-modal and long-horizon modeling, they are widely used in reinforcement learning (Janner et al., 2022; Ding et al., 2024, 2025), imitation learning (Wu et al., 2024b; Pearce et al., 2023), and motion planning (Reuss et al., 2023; Sridhar et al., 2023; Xian and Gkanatsios, 2023; Prasad et al., 2024; Saha et al., 2024). In robotics, DP (Chi et al., 2023) and DP3 (Ze et al., 2024) adopt diffusion model as the generative framework and condition the observations in the neural network to denoise the random noise into actions while flow based policies (Zhang et al., 2025; Hu et al., 2024) use the flow matching framework. Many other works aim to eliminate the dependence on random noise and generate actions directly from observations. BRIDGER (Chen et al., 2024) first trains a policy to obtain coarse actions and then refines them via a generative model. Another concurrent work, VITA (Gao et al., 2025), achieves noise-free generation by applying flow matching in the latent space of images and actions, after which actions are decoded from it. Our method models the multi-modal observations in the diffusion SDE trajectory via diffusion bridge. However, simply applying the diffusion bridge does not work with the modality and shape mismatch unresolved. Building upon the modality fusion and aligner makes sampling of BridgePolicy start from an informative prior instead of the uninformative noise and directly generate the executable actions, allowing the policy to effectively exploit observations and generate more precise and reliable behaviors.
@@ 2.2 Diffusion Bridges for Generative Modeling
Diffusion bridges enable the transition between two arbitrary distributions without the need to initiate the diffusion process from the random noise, which provides a more principled and mathematically exact generation paradigm for one-to-one distribution transformation. On one hand, Diffusion Schrödinger Bridges (Liu et al., 2023; Shi et al., 2024; De Bortoli et al., 2021b; Somnath et al., 2023; Chen et al., 2021; Deng et al., 2024; Wang et al., 2025) aim to determine a stochastic process by solving an entropy-regularized optimal transport problem between two distributions. However, its high computational complexity, particularly pronounced in high-dimensional settings or constraints, poses significant challenges for direct optimization. Other works (Zhou et al., 2023; Yue et al., 2023) incorporate Doob’s $h$-transform into forward SDE, delivering remarkable performance on image restoration benchmarks. UniDB (Zhu et al., 2025) reformulates diffusion bridges through an SOC-based optimization problem, proving that Doob’s $h$-transform is a special case of SOC theory, thereby unifying and generalizing existing $h$-transform-based diffusion bridge methods. In this work, we leverage the ability of diffusion bridge to connect heterogeneous distributions, specifically the multi-modal observation distribution and the action distribution of an expert policy, and harness this mechanism for policy learning.
@@ 3.1 BridgePolicy
Existing policy learning frameworks (Chi et al., 2023; Ze et al., 2024; Zhang et al., 2025; Hu et al., 2024) aim to learn a policy $\pi:\boldsymbol{O}\rightarrow\boldsymbol{A}$, which predict the actions $\boldsymbol{a}\in\boldsymbol{A}$ given the observations $\boldsymbol{o}\in\boldsymbol{O}$ which include 3D point cloud as the visual input and robot state and the actions are chunked into a short sequence of a trajectory finishing the task. Previous works such as DP3 (Ze et al., 2024) and FlowPolicy (Zhang et al., 2025) formulate the policy as a conditional generative model, where actions are generated from random noises under the condition of observations $\boldsymbol{o}$ (Chi et al., 2023; Ze et al., 2024; Zhang et al., 2025). They merely treated the informative observations as conditional signals to guide the neural network during denoising.
Forward Process of BridgePolicy. Motivated by the recent advances of Diffusion Bridge (Zhou et al., 2023; Yue et al., 2023; Zhu et al., 2025), where conditions can be explicitly modeled in the forward diffusion SDE trajectory, we directly integrate the observations into the diffusion process to generate a more controllable and precise robot action. Here, our BridgePolicy adopt a unified diffusion bridge framework (Zhu et al., 2025), formulated under the Stochastic Optimal Control (SOC) theory, as the core learning paradigm. To be consistent with the notations of previous diffusion for visuomotor policy learning (Chi et al., 2023), we denote $\boldsymbol{a}_{t}$ as the interpolated robot action state at time $t$. Assuming that the action $\boldsymbol{a}$ and the observation $\boldsymbol{o}$ share the same dimension, we define the endpoints of our forward diffusion bridge process (2) as follows:
@@
  • The terminal state $\boldsymbol{a}_{T}$: we set the terminal state to be the observation, i.e., $\boldsymbol{a}_{T}=\boldsymbol{o}$. It is a pivotal departure from standard diffusion-based or flow-based policy where $\boldsymbol{a}_{T}\sim\mathcal{N}(0,I)$ is typically the Gaussian noise.
According to UniDB (Zhu et al., 2025), we can construct the forward process of diffusion bridge as the following SOC optimization problem:
$\begin{gathered}\min_{\mathbf{u}_{t,\gamma}}\ \mathbb{E}\left[\int_{0}^{T}\frac{1}{2}\left\|\mathbf{u}_{t,\gamma}\right\|_{2}^{2}dt+\frac{\gamma}{2}\left\|\boldsymbol{a}_{T}^{u}-\boldsymbol{a}_{T}\right\|_{2}^{2}\right]\\ \text{s.t.}\ \mathrm{d}\boldsymbol{a}_{t}=[\theta_{t}(\boldsymbol{a}_{T}-\boldsymbol{a}_{t})+g_{t}\mathbf{u}_{t,\gamma}]\mathrm{d}t+g_{t}\mathrm{d}\boldsymbol{w}_{t},\ \boldsymbol{a}_{0}^{u}=\boldsymbol{a}_{0},\end{gathered}$ (1)
@@
  $\boldsymbol{z}_{obs}=\mathbf{softmax}({\boldsymbol{z}_{pc}\cdot\boldsymbol{z}_{s}^{T}}/{\sqrt{d_{s}}})\boldsymbol{z}_{s}$
@@
where we follow the $l_{1}$ norm in UniDB (Zhu et al., 2025) regarding the order of the norm, and provide the ablation study in Appendix. Consequently, BridgePolicy naturally integrates the informative observations $\boldsymbol{a}_{T}=\boldsymbol{o}$ both in the SDE trajectories and the neural networks and aims to learn a stochastic trajectory that transforms the observation $\boldsymbol{o}$ into a plausible action $\boldsymbol{a}$.
@@
 Initialize $\boldsymbol{a}_{T}=\boldsymbol{z}_{obs}=\mathbf{softmax}({\boldsymbol{z}_{pc}\cdot\boldsymbol{z}_{s}^{T}}/{\sqrt{d_{s}}})\boldsymbol{z}_{s}$
@@
Return Actions $\boldsymbol{a}$}
Sampling Algorithm of BridgePolicy. After we obtain the optimal model $\boldsymbol{a}^{*}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$, the decision making module of our BridgePolicy can be implemented as an iterative generation procedure starting from the given observation $\boldsymbol{a}_{T}=\boldsymbol{o}$. As for specially designed solvers, UniDB++ (Pan et al., 2025) provides a training-free acceleration algorithm with data prediction $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$ that directly estimates the fixed and smooth target $\boldsymbol{a}_{0}$. With some notations $\rho_{t}=e^{\bar{\theta}_{t}}(1-e^{-2\bar{\theta}_{t}})$, $\kappa_{t,\gamma}=e^{\bar{\theta}_{t:T}}((\gamma\lambda^{2})^{-1}+1-e^{-2\bar{\theta}_{t:T}})$, $c_{1}=(\gamma\lambda^{2})^{-1}e^{2\bar{\theta}_{T}}$, $c_{2}=e^{2\bar{\theta}_{T}}-1$, $D=2c_{1}c_{2}/(c_{1}+c_{2})^{3}$, $E=c_{2}^{2}/(c_{1}+c_{2})^{2}$, and $F=c_{1}^{2}/(c_{1}+c_{2})^{2}$ in UniDB++, the closed-form updating rule from time step $t$ to $s$ is formed as
$\begin{gathered}\begin{aligned} \boldsymbol{a}_{t}&=\frac{\kappa_{t,\gamma}\kappa_{s}\rho_{t}}{\kappa_{s,\gamma}\kappa_{t}\rho_{s}}\boldsymbol{a}_{s}+\left(\frac{\kappa_{t,\gamma}}{\kappa_{0,\gamma}}-\frac{\kappa_{t,\gamma}\rho_{t}\kappa_{s}}{\kappa_{0,\gamma}\kappa_{t}\rho_{s}}\right)\boldsymbol{a}_{\theta}(\boldsymbol{a}_{s},\boldsymbol{a}_{T},s)\\ &+\left(1-\frac{\kappa_{t,\gamma}\kappa_{s}\rho_{t}}{\kappa_{s,\gamma}\kappa_{t}\rho_{s}}+\frac{\kappa_{t,\gamma}\rho_{t}\kappa_{s}}{\kappa_{0,\gamma}\kappa_{t}\rho_{s}}-\frac{\kappa_{t,\gamma}}{\kappa_{0,\gamma}}\right)\boldsymbol{a}_{T}+\delta^{d}_{s:t}\boldsymbol{\epsilon},\end{aligned}\\ \begin{aligned} (\delta^{d}_{s:t})^{2}&=\frac{\lambda^{2}\kappa^{2}_{t,\gamma}\rho^{2}_{t}}{\kappa^{2}_{t}}\Bigg[E\frac{e^{2\bar{\theta}_{s}}-e^{2\bar{\theta}_{t}}}{(e^{2\bar{\theta}_{t}}-1)(e^{2\bar{\theta}_{s}}-1)}\\ &+D\log\frac{\kappa_{s,\gamma}\rho_{t}}{\kappa_{t,\gamma}\rho_{s}}-F(\frac{e^{-\bar{\theta}_{T}-\bar{\theta}_{t}}}{\kappa_{t,\gamma}}-\frac{e^{-\bar{\theta}_{T}-\bar{\theta}_{s}}}{\kappa_{s,\gamma}})\Bigg],\end{aligned}\end{gathered}$ (4)